OpenStudy (roadjester):

integral(f(x))dx from 0 to pi where f(x)= sinx if 0<=x

OpenStudy (anonymous):

-2

OpenStudy (roadjester):

not -2

OpenStudy (roadjester):

that's what i got too

OpenStudy (anonymous):

would you not just add the integral of sin(x) from 0 to pi/2 to the integral of cos(x) from pi/2 to pi? which is 1 + (-1) = 0, this is a piecewise defined function

OpenStudy (roadjester):

ok, the answer is 0, can u elaborate please?

OpenStudy (anonymous):

note the integral of sinx is -cosx

OpenStudy (roadjester):

ok, i can follow that

OpenStudy (roadjester):

i also know that the integral of cosx is sinx

OpenStudy (roadjester):

it's the evaluation idk how to do

OpenStudy (anonymous):

f(x) = sin(x) for 0< x </2 and f(x) = cos(x) for pi/2< x <pi so in evaluating the integral of sin(x), which is -cos(x) you get: -cos(pi/2) - (-cos(0)) which is 0 - (-1) which is 1. then evaluate the integral of cos(x) which is sin(x): sin(pi) - sin(pi/2) which is -1 - 0 = -1 add the values and you will get 0

OpenStudy (roadjester):

but aren't the 0, pi/2, and pi, restrictions of the graph? I see your logic but the question was integral of f(x)dx where f(x)=sin(x) if 0< = x < pi/2 and f(x)=cos(x) if pi/2< = x <= pi or is that irrelevant?

OpenStudy (anonymous):

so this says, f(x) = sin(x) where x is greater than or equal to x and less than pi/2, and f(x) = cos(x) where x is greater than or equal to pi/2 and less than or equal to pi. all values you mentioned are defined in the graph (0,pi/2,pi) you are finding area under curve from [0,pi/2) and [pi/2,pi]

OpenStudy (roadjester):

COOL! THANKS MAN (or girl/woman/miss, don't know your gender sorry) YOU ROCK!